OpenStudy (anonymous):

saifoo help me i just need to understand one thing saifoo help me i just need to understand one thing @Mathematics

OpenStudy (saifoo.khan):

Sure. =)

OpenStudy (karatechopper):

ima try 2! learn new thing everyday can i join?

OpenStudy (saifoo.khan):

Sure. $200 each.

OpenStudy (saifoo.khan):

plus $400 admission fee.

OpenStudy (anonymous):

ok when ur graphing an equation like for an example y-1=3(x+6) u use the slope intercept form which is 3 and -1 ? how will that be

OpenStudy (saifoo.khan):

y-1=3(x+6) Solve this till u get it in the form of y=mx+b y - 1 = 3x + 18 y = 3x + 18 -1 y = 3x + 17 So the slope = 3. y intercept = 17

OpenStudy (anonymous):

How come all of the bills has the same code

OpenStudy (saifoo.khan):

LOL, call 911 asap. KFC has her own printing press.

OpenStudy (karatechopper):

they r all 100 dollar bills so i had to pay 600 and i pasted it 6 times cuz 100 times 6 is 600

OpenStudy (anonymous):

but my teacher tought us another way to do it without computation

OpenStudy (karatechopper):

hahahhahahahahhaha LMFAO

OpenStudy (karatechopper):

hey i learned ur question 2-4 days ago i can help in a fast and easy way

OpenStudy (saifoo.khan):

the point which u are talking about will be like, (1,-6)

OpenStudy (saifoo.khan):

insert that in equation, x2-x1 = m (y2-y1)

OpenStudy (anonymous):

then that means the other people were wrong

OpenStudy (karatechopper):

so tht is the equation u go off of^

OpenStudy (saifoo.khan):

idk about others.. Lol.

OpenStudy (karatechopper):

so... ur equation: y-1=3(x+6)

OpenStudy (anonymous):

im alittle confused

OpenStudy (saifoo.khan):

where?

OpenStudy (anonymous):

ya thats my equation

OpenStudy (karatechopper):

tht equation can tell u ur points and u can put it into slope intercept form or u can solve it from ur equation which dp u prefer?

OpenStudy (anonymous):

my teachers formula is different

OpenStudy (karatechopper):

?

OpenStudy (karatechopper):

do u have to go by the teacher?

OpenStudy (anonymous):

into slope intercept form

OpenStudy (saifoo.khan):

slope intercept = y=mx+b

OpenStudy (saifoo.khan):

what's so confusing? o_O

OpenStudy (anonymous):

ya but how do i get the two points

OpenStudy (karatechopper):

ah ok i will help u solve it into tht y-1=3(x+6) y-1=3x-18 y-1+1=3x-18+1 y=3x-17: slope intercept form

OpenStudy (karatechopper):

the two points is easy

OpenStudy (karatechopper):

i can show u tht too

OpenStudy (saifoo.khan):

simply draw the line.. then get any two points.

OpenStudy (saifoo.khan):

KFC u did a mistake in signs.

OpenStudy (anonymous):

omg thats it thats what it was karate got it thats the way i was looking for

OpenStudy (karatechopper):

yay!

OpenStudy (karatechopper):

and how saif

OpenStudy (anonymous):

did he mess up on his signs?

OpenStudy (karatechopper):

o i see hold on wlbrt i got this and saif did too

OpenStudy (saifoo.khan):

i showed that above as well. LOL it's +17.

OpenStudy (anonymous):

oh

OpenStudy (karatechopper):

hold on i got this!

OpenStudy (saifoo.khan):

ah ok i will help u solve it into tht y-1=3(x+6) y-1=3x+18 y-1+1=3x+18+1 y=3x+19: slope intercept form

OpenStudy (saifoo.khan):

infact i did that incorrect as well. LOL.

OpenStudy (karatechopper):

y-1=3(x+6) y-1=3x+18 y-1+1=3x+18+1 y=3x+19: slope intercept form

OpenStudy (karatechopper):

haha o knew it!

OpenStudy (karatechopper):

u copied my thing saif!

OpenStudy (anonymous):

oh ya now explain how to graph it

OpenStudy (saifoo.khan):

Lol, i will pay u the rent for copying @kfc.

OpenStudy (karatechopper):

i can do tht!! can i do it! ooh! ooh!

OpenStudy (saifoo.khan):

proceed please.

OpenStudy (anonymous):

r u ther

OpenStudy (karatechopper):

k so y=3x+19 is what u have to graph 3 is ur slope cuz y=mx+b says m is slope so then u have the 19 and tht is ur y intercept. u go on the y axis and u go to positive 19 and plot. then if u have learned the rise over run formula then u would go up 3 and over one

OpenStudy (karatechopper):

cuz slope in a frac would b 3/1 or y/x

OpenStudy (anonymous):

oh ok

OpenStudy (karatechopper):

understand?

OpenStudy (anonymous):

u plot at 19

OpenStudy (anonymous):

right

OpenStudy (anonymous):

safoo

OpenStudy (karatechopper):

o and saif wheres my 700 dollars for plaijurism?

OpenStudy (anonymous):

hello

OpenStudy (anonymous):

kar

OpenStudy (saifoo.khan):

oh sorry, im back.

OpenStudy (saifoo.khan):

plagiarism*

OpenStudy (saifoo.khan):

so where are u stuck noW?

OpenStudy (karatechopper):

LOL i cant spell geez

OpenStudy (saifoo.khan):

LOL.

OpenStudy (karatechopper):

anybody see mimi?